a is bomb initially at rest when it suddenly explodes into 3 pieces. a 5kg piece goes northat 6m/s, and a 2kg piece goes east at 20m/s. The speed of the 3rd piece is 25m/s. what is the mass of the 3rd piece?
 

Grade:12th pass

1 Answers

Arun
6 years ago
Dear Student
 
conservation of momentum 

Let m be the unknown mass 
North is zero degrees and angles increase clockwise as on a map or compass. 

In the North direction 

5 kg(6 m/s) + m(25cosθ) = 0 
m(25cosθ) = -30 

In the East direction 

2 kg(20 m/s) + m(25sinθ) = 0 
m(25sinθ) = -40 

m(25sinθ) / m(25cosθ) = 40/30 
tanθ = 1.333 
θ = 233.13° (which is close to WSW) 

m(25cos233.13) = -30 
m = 2 kg 

or 

m(25sin233.13) = -40 
m = 2 kg 
 
Regards
Arun (askIITians forum expert)